OpenStudy (anonymous):

5.8 x 10^6 mg/s in grams per minute

OpenStudy (anonymous):

divide by 1000 multiply by 60

OpenStudy (anonymous):

?

OpenStudy (anonymous):

There are 1000 mg in a gram, and 60 seconds in a minute. To make your math easier, you could just count the zeros and subtract from the exponent. 5.8 * 10^6 becomes 5.8*10^3 (divide by 1000) then you multiply by 60(6*10) =5.8*6 *10^(3+1) =34.8*10^4 or 3.48 * 10^5.